OpenStudy (anonymous):

Use the Comparison Test to determine whether the series is convergent of divergent. sum of sqrt(n)/(n-1) from n=2 to infinite

OpenStudy (anonymous):

\[\sum_{n=2}^{\infty} \frac{ \sqrt{n} }{ n-1 }\]

OpenStudy (anonymous):

evaluate some values of n (ex: n = 1, 2, 3, 4, ...) If, when evaluated, the values seems to be approaching a certain number, then it is convergent. if not, then it is diveregent